OpenStudy (trwatkins1):

What is the solution to the equation? b + 0.8 = –1 b =

OpenStudy (anonymous):

-1.8

OpenStudy (kittiwitti1):

Isolate the b.\[b+0.8=-1\rightarrow b+0.8-0.8=1-0.8\rightarrow b=?\]

OpenStudy (kittiwitti1):

@Cat_LadyxD Please do not give direct answers, that is against the OpenStudy Code of Conduct.